1.The healthcare provider orders 240 mg of a drug IVPB q6h. The drug is to be is administered over 30 minutes and should be diluted in at least 5 mL of fluid. The​ child's IV is to infuse at 50 mL per hour. After adding the medication to the​ VCC, you will add IV fluid to what​ level?

2. ​Order: NS​ 1,000 mL IV stat at 125​ mL/h for 8 hours. After 4​ hours, 600 mL remain to be infused. The infusion is behind schedule. Compute the new flow rate in​ mL/h so that the infusion will finish on time. Note that flow rates may not be adjusted by more than​ 25% of the original rate. To what rate would you reset the drip​ rate, or would you contact the​ prescriber?

3. A drug is ordered at 10​ mg/kg PO daily in two divided doses. The label on the bottle​ reads, "25​ mg/tab." Calculate the number of tablets you should administer per dose for a child who weighs 22 pounds.
Answer #1

3. Ordered drug= 10​ mg/kg PO daily in two divided doses.

Weight of the child= weighs 22 pounds= 22/2.2kg= 10 kg weight

As per weight of the child required dose= 10 x10 = 100 mg

The label on the bottle​ reads, "25​ mg/tab."

required tablets for the child= 100/25= 4 tablets daily